skip to content
 
Suzaku Guest Observer Facility

In-Orbit Status & Updates

Last Updated 2005 August 30


Significant Events

  • Aug 18: HXD now operational - see the first light spectrum
  • Aug 12-13: XIS now operational - see the first light images
  • Aug 9: Loss of all liquid helium from the XRS
  • Jul 23-29: Attitude Control System is being tuned. Star trackers are now used to calibrate the gyro data. Suzaku is approaching routine attitude operation, although it is not quite there yet.
    • XRS status: ADR was opeated, and the detector achieved the operational temperature of 60 mK. The calibration source data already show an energy resolution of 7 eV, and further tuning of parameters is being attempted to improve this. Filter wheel operation was tested successfully.
    • XIS status: the thermo-electric coolers work well, and the CCD chips achieved -90C as intended. Cosmic ray and calibration source events are already seen in all 4 detectlors. Various observation modes (5x5, 3x3 etc.) have been tested successfully.
    • HXD status: Various electronics boxes have been switched on. High voltage have been successfully applied to the phototubes and PIN diodes. All 46 phototubes and 64 photodiodes are confirmed operational.
  • Jul 20-22: After a series of thruster burns, Suzaku has achieved an altitude of 568 km and an orbital eccentricity of 0.0002. The set-up of the star tracker and gyros are also complete, and Suzaku has transitioned to normal attitude operation.
    • XRS status: Mechanical cooler power has been increased to its maximum value (34.7W). Ne temperature has reached an equilibrium value of just under 16 K, and He is approaching equilibrium around 1.17 K. The Adiabatic Demagnetization Refregirator (ADR) is being tested in preparation for the first full gross cycle control (GCC) scheduled for next week (this will cool the microcalorimeter to its operational temperature of 60 mK).
  • Jul 15-19: Suzaku was successfully brought out of the safe hold. The star trackers were turned on. Attitude maneuvers were tested, then a major maneuver was carried out in several steps to orient the spacecraft for the delta-V burns. Several delta-V burns were performed during this period, successfully raising the perigee; there will be a few more delta-V burns over the next several days.
    • XRS status: various set-up commands are uploaded as time permits.
  • Jul 13: Suzaku has been placed in a planned safe hold.
    • XRS Status: All XRS pixels are confirmed to be working electrically. The mechanical cooler is operating successfully.
  • Jul 12: The extensible optical bench (EOB) was successfully extended. The pulse counter of the deploy mechanism recorded 77 pulses, exactly as expected. Confirming signals from all four latch mechanisms were received at 03:33:40 UT during the 4th contact at USC.
    • XRS status: CAP setup commands executed, including the turning-on of the Junction Field effect Transistor (JFET) amplifiers. All JFETs are operational, and the anti-coincidence detector is working well.
  • Jul 11: Satellite was de-spinned, and 3-axis attitude control was established. Solar panels were deployed. Power control system is working well and stable.
    • XRS status: the ADR Control and Houskeeping Electronics (ACHE) box is on, and the data look nominal. The dewar temperatures are properly decreasing. All XRS valves that were commanded to open did so without any problems. The mechanical cooler, the Calorimeter Analog Processor (CAP) and the Calorimeter Digital Processor (CDP) are on and working properly.
  • Jul 10: Launch!


If you have any questions concerning Suzaku, visit the Feedback form.

This file was last modified on Monday, 14-Aug-2006 13:41:37 EDT

NASA Astrophysics

  • FAQ/Comments/Feedback
  • Education Resources
  • Download Adobe Acrobat
  • A service of the Astrophysics Science Division (ASD) at NASA/ GSFC

    Suzaku Project Scientist: Dr. Robert Petre
    Responsible NASA Official: Phil Newman

    Privacy Policy and Important Notices.